OpenStudy (anonymous):

Solve. 6(a + 3) = 21 – 3(1 – 2a) A. no solution B. all real numbers C. –3 D. –1

OpenStudy (anonymous):

distribute then then combine like terms

OpenStudy (anonymous):

so the answer is C ?

OpenStudy (anonymous):

hurry :(

OpenStudy (anonymous):

6a+18=21 - 3- 6a 6a+6a+18=21-3 12a+18=21-2 12a=21-18-3 12a=3-3 12a=0 a=0 its no solution

OpenStudy (anonymous):

thank you.